3. Four point charges are placed at the four corners of a square that is 30 cm on each side. Two of the 4 charges are +2μC and two are -2μC. Find the potential at the center of the square.

 

4. The potential difference between two parallel plates of a condenser that are 4 mm apart is 800 V. What is the potential gradient?


1
Expert's answer
2022-03-27T18:21:00-0400

3. The net potential

"V_{net}=k\frac{q_1}{r_1}+k\frac{q_2}{r_2}+k\frac{q_3}{r_3}+k\frac{q_4}{r_4}""r_1=r_2=r_3=r_4=\frac{\sqrt{2}}{2}a\\
=\frac{\sqrt{2}}{2}*0.3=0.21\:\rm m"

"V_{net}=\frac{9*10^9}{0.21}(2+2-2-2)*10^{-6}=0\:\rm V"

4.

"\frac{\Delta V}{\Delta d}=\frac{800\:\rm V}{0.004\: \rm m}=2*10^5\:\rm V/m"